UNDULATING POTENTIAL ENERGY SURFACES FOR THE RYDBERG STATES OF SMALL MOLECULES CONTAINING A METAL ATOM

N. GEUM, Department of Chemistry, Dankook University, Cheonan, 330-714, South Korea; \underline{G.-H. JEUNG}, Laboratoire Aimé Cotton (CNRS UPR3321) and ASCI (CNRS UPR9029, Bât. 505-506, Campus d’Orsay, 91405 Orsay, France.